### v2

# Uniform counterexamples to the convergence problem for periodic dispersive equations with a polynomial symbol

Abstract.
In the setting of Carleson’s convergence problem for the fractional Schrödinger equation $i\,\partial_{t}u+(-\Delta)^{a/2}u=0$ with $a>1$ in $\mathbb{R}^{d}$, which has Fourier symbol $P(\xi)=|\xi|^{a}$, the Sobolev exponent $d/(2(d+1))$ is sufficient, but it is not known whether this condition is necessary. In this article, we show that in the periodic problem in $\mathbb{T}^{d}$ the exponent $d/(2(d+1))$ is necessary for all non-singular polynomial symbols $P$ regardless of the degree of $P$. Among the differential operators covered, we highlight the natural powers of the Laplacian $\Delta^{k}$ for $k\in\mathbb{N}$.
1. Introduction
For a function $P:\mathbb{R}^{d}\to\mathbb{R}$ and $D=(\partial_{1},\ldots,\partial_{d})$, consider the periodic initial value problem
$\left\{\begin{array}[]{ll}i\partial_{t}u+P(D)u=0,&x\in\mathbb{T}^{d},\quad t\in\mathbb{R},\\ u(x,0)=f(x),&x\in\mathbb{T}^{d},\end{array}\right.$ (1.1)
whose solution we denote by
$e^{itP(D)}f(x)=\sum_{n\in\mathbb{Z}^{d}}\widehat{f}_{n}\,e^{2\pi i(nx+P(n)t)}.$ (1.2)
We study Carleson’s problem of convergence to the initial datum: for which $s$ do we get
$\lim_{t\to 0}e^{itP(D)}f(x)=f(x)\qquad\text{ for almost every }x\in\mathbb{T}^{d},\qquad\forall f\in H^{s}(\mathbb{T}^{d})\quad\text{?}$ (1.3)
It is typical to encode this problem using the critical exponent
$s_{P}(\mathbb{T}^{d})=\inf\big{\{}\,s\,:\,\lim_{t\to 0}e^{itP(D)}f(x)=f(x)\quad\text{a.e. }x\in\mathbb{T}^{d},\quad\forall f\in H^{s}(\mathbb{T}^{d})\,\big{\}}.$ (1.4)
Replacing $\mathbb{T}^{d}$ by $\mathbb{R}^{d}$, and the Fourier series (1.2) by the corresponding Fourier integral, one gets the analogous Euclidean problem with the critical exponent $s_{P}(\mathbb{R}^{d})$.
When $P(\xi)=|\xi|^{2}$, equation (1.1) is the Schrödinger equation, for which we now know [4, 9, 3, 11, 12] that
$s_{|\xi|^{2}}(\mathbb{R}^{d})=\frac{d}{2(d+1)}.$ (1.5)
The problem is open on $\mathbb{T}^{d}$, the best results [18, Prop 1 in Section 2], [8, Prop. 3.1 and 3.2] being
$\frac{d}{2(d+1)}\leq s_{|\xi|^{2}}(\mathbb{T}^{d})\leq\frac{d}{d+2}.$ (1.6)
A natural generalization to the quadratic symbol $|\xi|^{2}$ is $P(\xi)=|\xi|^{a}$ for $a>0$, which replaces the Laplacian operator $|D|^{2}=\Delta$ in (1.1) by its fractional powers $(-\Delta)^{a/2}$. We know that
$s_{|\xi|^{a}}(\mathbb{R}^{d})\leq\frac{d}{2(d+1)},\qquad\text{ for every }a>1.$ (1.7)
This was shown by Sjölin [19] when $d=1$. The case $d=2$ was proved by Cho and Ko [5] adapting the work by Du, Guth and Li [11]. In the same work [5, Remark 3.12], Cho and Ko assert that the result holds for $d\geq 3$ by adapting their proof following the methods of Du and Zhang [12]. This was also proved by Li, Li and Xiao [17]. The case $a\leq 1$ is considerably different and we refer ot [6] for a recent account of the state of the art.
Sjölin [19] also proved that (1.7) is sharp when $d=1$. However, counterexamples are scarce in $d\geq 2$. One of the main obstructions for $|\xi|^{a}$ is the loss of a simple algebraic structure [^1] Compare $\xi_{1}^{a}$ in $d=1$ and $(\xi_{1}^{2}+\xi_{2}^{2})^{a/2}$ in $d=2$, which has a more complicated structure especially when $a\not\in 2\mathbb{N}$.. For example, Bourgain’s proof [3] of the lower bound in (1.5) critically exploits the separation of variables of the symbol $|\xi|^{2}=\xi_{1}^{2}+\ldots+\xi_{d}^{2}$, and does not naturally adapt to $|\xi|^{a}$. To avoid these algebraic obstructions, An, Chu and Pierce [1] adapted Bourgain’s construction for the separated symbol
$\xi_{1}^{k}+\xi_{2}^{k}+\ldots+\xi_{d}^{k},\qquad\qquad k\in\mathbb{N}\setminus\{1\},$ (1.8)
which corresponds to the differential operator $\partial_{1}^{k}+\partial_{2}^{k}+\ldots+\partial_{d}^{k}$. Using Deligne’s theorem to estimate Weyl sums of order $k$, they proved
$s_{\xi_{1}^{k}+\ldots+\xi_{d}^{k}}(\mathbb{R}^{d})\geq\frac{1}{4}+\frac{d-1}{4((k-1)d+1)}.$ (1.9)
The same results hold for polynomial symbols of degree $k$ whose homogeneous part has the form
[…]
as shown by the first author and Ponce-Vanegas [15]. Chu and Pierce [7] extended these results to a broader family of symbols, classifying them with an intertwining rank $r$ that quantifies the amount of cross-terms in the symbol, giving a non-trivial lower bound that depends on $k$ and $r$.
The symbols (1.8) can be regarded as an intermediate step towards the more natural $|\xi|^{2k}$, which represent the natural powers of the Laplacian $\Delta^{k}$ in the dispersive equations $i\partial_{t}u+\Delta^{k}u=0$. These symbols are polynomials so they preserve a favorable algebraic structure, but they do not separate variables if $k\geq 2$. In this note, we show that in the periodic setting this problem is irrelevant, obtaining a lower bound for the critical exponent that is independent of $k$.
Theorem 1.1.
Let $P\in\mathbb{Z}[X_{1},\ldots,X_{d}]$ be a polynomial with arbitrary $\deg P\geq 2$. Suppose that $P$ is non-singular in the precise way described in Theorem 4.1. Then,
$s_{P}(\mathbb{T}^{d})\geq\frac{d}{2(d+1)}.$ (1.11)
In particular, the result holds for:
  • $P(\xi)=|\xi|^{2k}$ for any $k\in\mathbb{N}$, which corresponds to the powers $\Delta^{k}$ of the Laplacian.
  • $\xi_{1}^{k}+\ldots+\xi_{d}^{k}$ for any $k\in\mathbb{N}\setminus\{1\}$.
This result is in contrast with the lower bound (1.9) for $\mathbb{R}^{d}$, which depends on $k$. Moreover,
$\frac{1}{4}+\frac{1}{4(d(k-1)+1)}<\frac{d}{2(d+1)},\qquad\forall k>2.$ (1.12)
Together with the results known for $\mathbb{T}^{d}$ and $\mathbb{R}^{d}$, Theorem 1.1 suggest that the correct Sobolev exponent may indeed be $d/(2(d+1))$ for all $a>1$. This, of course, remains to be confirmed.
The proof of the theorem relies in the well-known characterization of the almost everywhere convergence in terms of $H^{s}\to L^{p}$ maximal estimates for the operator $e^{itP(D)}$. In this case, we will give a counterexample to the maximal estimate
$\big{\lVert}\sup_{t>0}|e^{itP(D)}f|\big{\rVert}_{L^{2}(B(0,1))}\lesssim\lVert f\rVert_{s},\qquad\forall f\in H^{s}(\mathbb{T}^{d})$ (1.13)
when $s<d/(2(d+1))$. This counterexample is inspired in the work by the first author and Lucà [13], and of Compaan, Lucà and Staffilani [8], and we use the techniques developed by An, Chu and Pierce in [1] and the first author and Ponce-Vanegas in [15].
A conjecture for the fractal problem
The fractal problem consists in strengthening the convergence condition to be satisfied almost everywhere with respect to some given Hausdorff measure. Let $\alpha\in[0,d]$ and let $\mathcal{H}^{\alpha}$ denote the $\alpha$-Hausdorff measure. The objective is to determine
$s_{P}(\mathbb{T}^{d},\alpha)=\inf\{\,s\,:\,\lim_{t\to 0}e^{itP(D)}f(x)=f(x)\quad\mathcal{H}^{\alpha}\text{-a.e. }x\in\mathbb{T}^{d},\quad\forall f\in H^{s}(\mathbb{T}^{d})\,\},$ (1.14)
with the analogous definition for the exponent $s_{P}(\mathbb{R}^{d},\alpha)$ in the Euclidean space $\mathbb{R}^{d}$. Of course, when $\alpha=d$ we recover the exponent $s_{P}(\mathbb{T}^{d})$ defined in (1.4).
For the classical symbol $|\xi|^{2}$, the Euclidean problem is solved when $\alpha\leq d/2$, the exponent being $s_{|\xi|^{2}}(\mathbb{R}^{d},\alpha)=(d-\alpha)/2$ [2, 20]. However, it is open when $\alpha>d/2$, the best results being
$s_{\text{sawtooth}}(\alpha)\leq s_{|\xi|^{2}}(\mathbb{R}^{d},\alpha)\leq\frac{d}{2(d+1)}(d+1-\alpha),\qquad\text{ for }\frac{d}{2}<\alpha<d.$ (1.15)
where $s_{\text{sawtooth}}(\alpha)$ is given in [14] and the upper bound in [12]. For the periodic problem, the current best results [13] are
$\frac{d}{2(d+1)}(d+1-\alpha)\leq s_{|\xi|^{2}}(\mathbb{T}^{d},\alpha)\leq\frac{d}{2(d+2)}(d+2-\alpha),\qquad\text{ for }\alpha\leq d.$ (1.16)
For the symbols $|\xi|^{a}$, the upper bound in (1.15) holds for $s_{|\xi|^{a}}(\mathbb{R}^{d},\alpha)$ for all $a>1$ [17]. The same heuristics that give Theorem 1.1 suggest that in the periodic case, the lower bound in (1.16) holds for any non-singular polynomials of higher degree.
Conjecture 1.2.
Let $P$ be a polynomial with the same hypothesis as in Theorem 1.1. Then,
$s_{P}(\mathbb{T}^{d},\alpha)\geq\frac{d}{2(d+1)}(d+1-\alpha),\qquad\text{ for }\alpha\leq d.$ (1.17)
The critical difference with respect to the case $\alpha=d$ is that the equivalence to the $H^{s}\to L^{2}$ maximal estimate for the Schrödinger operator need not be true in the fractal case $\alpha<d$. Thus counterexamples have to be built directly for the convergence problem. In particular, this implies computing the Hausdorff dimension of the sets of divergence, which is in general a hard problem. We aim to come back to this in the future.
2. The setup
Let $\psi:\mathbb{R}\to\mathbb{R}$ be a smooth function with support in $[1/4,2]$ such that $\psi(\xi)=1$ in $\xi\in[1/2,1]$. Define $\varphi:\mathbb{R}^{d}\to\mathbb{R}$ by
$\varphi(\xi)=\psi(\xi_{1})\,\ldots\,\psi(\xi_{d}),\qquad\xi=(\xi_{1},\ldots,\xi_{d})\in\mathbb{R}^{d}.$ (2.1)
Let $N\gg 1$ and define the datum
$f_{N}(x)=\sum_{n\in\mathbb{Z}^{d}}\varphi\Big{(}\frac{n}{N}\Big{)}e^{2\pi in\cdot x},$ (2.2)
where $n=(n_{1},\ldots,n_{d})\in\mathbb{Z}^{d}$ and $x=(x_{1},\ldots,x_{d})\in\mathbb{T}^{d}$. Direct computation shows that
$\lVert f_{N}\rVert_{H^{s}}^{2}=\sum_{n\in\mathbb{Z}^{d}}(1+|n|^{2})^{s}\,\varphi\Big{(}\frac{n}{N}\Big{)}^{2}\simeq N^{2s+d},$ (2.3)
while the corresponding solution has the form
$e^{itP(D)}f_{N}(x)=\sum_{n\in\mathbb{Z}^{d}}\varphi\Big{(}\frac{n}{N}\Big{)}e^{2\pi i(n\cdot x+P(n)t)}.$ (2.4)
3. Heuristics and choice of parameters
We first show the key ideas of the proof, avoiding technical details. The rigorous proof will be given in Section 4.
3.1. The solution
Following Bourgain’s idea in [3], choose rational space and time
$x=\frac{b}{q}\qquad\text{ and }\qquad t=\frac{1}{q},$ (3.1)
with $q\in\mathbb{N}$ and $b\in\mathbb{\{}0,1,\ldots,q-1\}^{d}$, so that we can turn the solution (2.4) into a Weyl sum
$e^{itP(D)}f_{N}(x)=\sum_{n\in\mathbb{Z}^{d}}\varphi\Big{(}\frac{n}{N}\Big{)}e^{2\pi i\frac{b\cdot n+P(n)}{q}}.$ (3.2)
For the sake of simplicity, assume for the moment that $P(\xi)=\xi_{1}^{k}+\ldots+\xi_{d}^{k}$. Assume also $q\ll N$. Counting cycles of length $q$, the solution splits into
$|e^{itP(D)}f_{N}(x)|\simeq\Big{|}\sum_{n\in\mathbb{Z},\,n\simeq N}e^{2\pi i\frac{bn+n^{k}}{q}}\Big{|}^{d}\simeq\Big{|}\frac{N}{q}\sum_{n=1}^{q}e^{2\pi i\frac{bn+n^{k}}{q}}\Big{|}^{d},$ (3.3)
where we slightly abuse notation by denoting each coordinate $b_{i}$ by $b$. If $k=2$, the sum is a quadratic Gauss sum which, up to choosing the correct $b$, is known to have size $\sqrt{q}$. Similarly, if $k\geq 3$, it is a consequence of Deligne’s theorem on the Weil conjectures that if $q$ is a prime, the sum in (3.3) also has size $\sqrt{q}$ for a positive proportion of $b\in\{0,1,\ldots,q-1\}$. Hence, we expect
$|e^{itP(D)}f_{N}(x)|\simeq\Big{(}\frac{N}{\sqrt{q}}\Big{)}^{d}.$ (3.4)
Estimate (3.4) is also valid for any non-singular polynomial $P$ of degree $k$, since Deligne’s theorem implies in general that
$\Big{|}\sum_{n_{1}=1}^{q}\cdots\sum_{n_{d}=1}^{q}e^{2\pi i\frac{b\cdot n+P(n)}{q}}\Big{|}^{d}\simeq q^{d/2}.$ (3.5)
We refer to Theorem 4.1 and the following discussion for the detailed statements and estimates.
3.2. The set of divergence
The estimate (3.4) holds as long as
$x=\frac{b}{q}\qquad\text{ and }\qquad t=\frac{1}{q}\qquad\text{ for }q\text{ prime and most }b\in\{0,1,\ldots,q-1\}^{d},$ (3.6)
where most $b$ means, say, at least half of them. However, the set of all such $x$ has, of course, zero measure. To gain volume, let $\delta\in\mathbb{R}^{d}$ and let
$x=\frac{b}{q}+\delta.$ (3.7)
This generates an additional $e^{2\pi in\cdot\delta}$ term in (3.2), which will be negligible if $|n\cdot\delta|\simeq N|\delta|<1$. Hence (3.4) holds even if $x\in B(b/q,1/N)$.
On the other hand, observe that the relevant information about $q$ in the estimate (3.4) is its size rather than its specific value. Hence, denote by $Q$ a new parameter such that $1\ll Q\ll N$, and let $\mathcal{P}_{Q}=\{q\simeq Q\,:\,q\text{ prime }\,\}$. Define
$X_{N}=\bigcup_{q\in\mathcal{P}_{Q}}\bigcup_{\text{ most }b\in\{0,1,\ldots,q-1\}^{d}}B\Big{(}\frac{b}{q},\,\frac{1}{N}\Big{)}.$ (3.8)
The discussion in Section 3.1 implies that
$x\in X_{N}\quad\Longrightarrow\quad\exists\,t\simeq\frac{1}{Q}\quad\text{ suchthat }\quad|e^{itP(D)}f_{N}(x)|\simeq\Big{(}\frac{N}{\sqrt{q}}\Big{)}^{d}.$ (3.9)
3.3. The measure of the set of divergence
Since $X_{N}\subset[0,1]^{d}$, we trivially have $|X_{N}|\leq 1$. We want to choose the smallest possible $Q$ that makes $|X_{N}|\simeq 1$. We give two arguments that complement each other to find such $Q$.
  1. 1.
    Dirichlet approximation. For any $x\in[0,1]^{d}$, there are infinitely many $q\in\mathbb{N}$ and $b\in\mathbb{Z}^{d}$ such that
    $\Big{|}x_{i}-\frac{b_{i}}{q}\Big{|}\leq\frac{1}{q^{(d+1/d)}},\qquad\text{ for all }i=1,\ldots,d.$ (3.10)
    This amounts to say that balls $B(b/q,1/q^{(d+1)/d})$ should suffice to cover the unit interval $[0,1]^{d}$. Hence, in view of (3.8), by choosing
    $N\simeq Q^{(d+1)/d}$ (3.11)
    it is natural to expect that $X_{N}$ covers $[0,1]^{d}$ in an efficient way.
  2. 2.
    Disjointness. That $X_{N}$ has good covering properties can be rephrased as the balls in the definition of $X_{N}$ not overlapping too much. If we assume that the balls are actually disjoint, then
    $|X_{N}|=\sum_{q\in\mathcal{P}_{Q}}\sum_{\text{most }b\,\in\,\{0,1,\ldots,q-1\}^{d}}\frac{1}{N^{d}}\simeq\frac{Q}{\log Q}\,Q^{d}\,\frac{1}{N^{d}}=\frac{1}{\log Q}\,\frac{Q^{d+1}}{N^{d}},$ (3.12)
    since $|\mathcal{P}_{Q}|\simeq Q/\log Q$. We will see in Section 4 that if we choose [^2] If we ignore the $\log Q$ term due to $q$ being prime, the choice (3.11) is precisely the smallest $Q$ that makes $|X_{N}|\simeq 1$. $Q$ as in (3.11), then it is indeed the case that
    $|X_{N}|\simeq\frac{1}{\log Q}.$ (3.13)
3.4. Disproving the maximal estimate
Gathering the estimates (2.3), (3.4) and (3.13), and the choice of the parameter $Q$ in (3.11), we immediately get
$\frac{\lVert\sup_{t}\big{|}e^{itP(D)}f_{N}\big{|}\rVert_{L^{2}}}{\lVert f_{N}\rVert_{H^{s}}}\gtrsim\Big{(}\frac{N}{\sqrt{Q}}\Big{)}^{d}\,N^{-s-d/2}\,|X_{N}|^{1/2}\simeq\frac{1}{(\log N)^{1/2}}\,N^{\frac{d}{2(d+1)}-s},$ (3.14)
which disproves the maximal estimate (1.13) if $s<\frac{d}{2(d+1)}$.
4. Rigorous proof of Theorem 1.1
In this section, we rigorously prove the heuristics outlined in Section 3. According to (3.4), we will work with $Q=N^{d/(d+1)}$, so in particular $1\ll Q\ll N$.
4.1. A lower bound for the solution
Let $b\in\mathbb{Z}^{d}$ and let $q\in\mathbb{N}$ be prime [^3] The estimate for the Weil sum in Theorem 4.1 is the only place where we need $q$ to be prime.. Let $\delta\in\mathbb{R}^{d}$. Define
$x=\frac{b}{q}+\delta,\qquad t=\frac{1}{q},\qquad\text{ with }\quad|\delta|\leq\frac{c}{N},$ (4.1)
with $c>0$ small enough. Then,
$e^{itP(D)}f_{N}(x)=\sum_{n\in\mathbb{Z}^{d}}\varphi\Big{(}\frac{n}{N}\Big{)}e^{2\pi i\frac{b\cdot n+P(n)}{q}}\,e^{2\pi i\delta\cdot n}=\sum_{n\in\mathbb{Z}^{d}}\zeta(n)\,e^{2\pi i\frac{b\cdot n+P(n)}{q}},$ (4.2)
where
$\zeta(n)=\varphi(n/N)e^{2\pi i\delta\cdot n}.$ (4.3)
We split the sum modulo $q$. Write $n=mq+r$ with $m\in\mathbb{Z}^{d}$ and $r\in\mathbb{F}_{q}^{d}$, where by $\mathbb{F}_{q}$ we denote the field of integers modulo $q$. By periodicity, we have
$e^{itP(D)}f_{N}(x)=\sum_{r\in\mathbb{F}_{q}^{d}}\Big{(}\sum_{m\in\mathbb{Z}^{d}}\zeta(mq+r)\Big{)}e^{2\pi i\frac{b\cdot r+P(r)}{q}}=\sum_{r\in\mathbb{F}_{q}^{d}}Z(r)e^{2\pi i\frac{b\cdot r+P(r)}{q}},$ (4.4)
where we define
$Z(r)=\sum_{m\in\mathbb{Z}^{d}}\zeta(mq+r),\qquad\qquad\text{ for }r\in\mathbb{F}_{q}^{d}.$ (4.5)
Decompose this function into frequencies by the discrete Fourier transform
$Z(r)=\sum_{\ell\in\mathbb{F}_{q}^{d}}\widehat{Z}(\ell)\,e^{2\pi i\frac{r\cdot\ell}{q}},\qquad\text{ with }\qquad\widehat{Z}(\ell)=\frac{1}{q^{d}}\,\sum_{r\in\mathbb{F}_{q}^{d}}Z(r)\,e^{-2\pi i\frac{r\cdot\ell}{q}},$ (4.6)
so that from (4.4) we get a main component and an error term,
$\begin{split}e^{itP(D)}f_{N}(x)&=\widehat{Z}(0)\,\sum_{r\in\mathbb{F}_{q}^{d}}e^{2\pi i\frac{b\cdot r+P(r)}{q}}+\sum_{r\in\mathbb{F}_{q}^{d}}\sum_{\ell\in\mathbb{F}_{q}^{d}\setminus\{0\}}\widehat{Z}(\ell)\,e^{2\pi i\frac{(b+\ell)\cdot r+P(r)}{q}}\\ &=M+E.\end{split}$ (4.7)
Main term
From the definitions of $Z$ and $\zeta$ in (4.5), direct computation shows
$\widehat{Z}(0)=\frac{1}{q^{d}}\,\sum_{r\in\mathbb{F}_{q}^{d}}Z(r)=\frac{1}{q^{d}}\,\sum_{r\in\mathbb{F}_{q}^{d}}\sum_{m\in\mathbb{Z}^{d}}\zeta(mq+r)=\frac{1}{q^{d}}\,\sum_{n\in\mathbb{Z}^{d}}\varphi\Big{(}\frac{n}{N}\Big{)}\,e^{2\pi i\delta\cdot n}\simeq\frac{N^{d}}{q^{d}}.$ (4.8)
To estimate the exponential sum, we have the following theorem which follows from a theorem of Deligne (see [10, Theorem 8.4] and [16, Theorem 11.43]) and Hilbert’s Nullstellensatz as explained in [15, Corollary 3.2].
Theorem 4.1.
Let $P\in\mathbb{Z}[X_{1},\ldots,X_{d}]$ such that $\deg P=k$. Let $P_{k}$ be the homogeneous part of $P$. Suppose that $\nabla P_{k}(x)\neq 0$ for all $x\in\mathbb{C}^{d}\setminus\{0\}$. Then, for every $q$ prime such that $q\nmid k$,
$\Big{|}\sum_{n\in\mathbb{F}_{q}^{d}}e^{2\pi i\frac{P(n)}{q}}\Big{|}\leq(k-1)q^{d/2}.$ (4.9)
From the upper bound in Theorem 4.1 we can get a lower bound for the exponential sum in $M$ as follows. Call
$S(b)=\sum_{r\in\mathbb{F}^{d}_{q}}e^{2\pi i\frac{P(r)+b\cdot r}{q}},\qquad b\in\mathbb{F}_{q}^{d}.$ (4.10)
Then,
$\sum_{b\in\mathbb{F}_{q}^{d}}|S(b)|^{2}=\sum_{r,s\in\mathbb{F}_{q}^{d}}\Big{(}e^{2\pi i\frac{P(r)-P(s)}{q}}\,\sum_{b\in\mathbb{F}_{q}^{d}}e^{2\pi i\frac{r-s}{q}\cdot b}\Big{)}=\sum_{r\in\mathbb{F}_{q}^{d}}q^{d}=q^{2d},$ (4.11)
because the inner sum is zero unless $r=s$. Let $c>0$ and define $G(q)=\{\,b\in\mathbb{F}_{q}^{d}\,:\,|S(b)|\geq c\,q^{d/2}\,\}$. By Theorem 4.1 we have $|S(b)|\leq(k-1)q^{d/2}$ for all $b$, so
$q^{2d}=\sum_{b\in\mathbb{F}_{q}^{d}}|S(b)|^{2}\leq|G(q)|\,(k-1)^{2}q^{d}+q^{d}\,(cq^{d/2})^{2},$ (4.12)
which implies
$|G(q)|\geq q^{d}\,\frac{1-c^{2}}{(k-1)^{2}}\simeq q^{d}.$ (4.13)
This shows that
$\exists G(q)\subset\mathbb{F}_{q}^{d}\quad\text{ with }\quad|G(q)|\gtrsim q^{d}\qquad\text{ such that }\quad\Big{|}\sum_{r\in\mathbb{F}^{d}_{q}}e^{2\pi i\frac{P(r)+b\cdot r}{q}}\Big{|}\gtrsim q^{d/2}\quad\text{ for all }b\in G(q).$ (4.14)
Joining this with (4.8), we get
$M=|\widehat{Z}(0)|\,\Big{|}\sum_{r\in\mathbb{F}_{q}^{d}}e^{2\pi i\frac{b\cdot r+P(r)}{q}}\Big{|}\gtrsim\frac{N^{d}}{q^{d}}\,q^{d/2},\qquad\forall b\in G(q).$ (4.15)
Error term
Using the Weil bound in Theorem 4.1, we first bound the error term by
$|E|\leq\Bigg{(}\sup_{\ell\in\mathbb{F}_{q}^{d}}\Big{|}\sum_{r\in\mathbb{F}_{q}^{d}}\,e^{2\pi i\frac{(b+\ell)\cdot r+P(r)}{q}}\Big{|}\Bigg{)}\,\sum_{\ell\in\mathbb{F}_{q}^{d}\setminus\{0\}}\big{|}\widehat{Z}(\ell)\big{|}\,\,\lesssim\,\,q^{d/2}\sum_{\ell\in\mathbb{F}_{q}^{d}\setminus\{0\}}\big{|}\widehat{Z}(\ell)\big{|}.$ (4.16)
To exploit the decay of the Fourier coefficients $\widehat{Z}(\ell)$, we sum by parts as follows. Denote the canonical basis by $\{e_{j}\}_{j=1}^{d}$ and define the discrete second derivatives by the differences
$\Delta_{j}g(n)=g(n+e_{j})-2g(n)+g(n-e_{j}),\qquad n=(n_{1},\ldots,n_{d})\in\mathbb{Z}^{d},$ (4.17)
for any function $g:\mathbb{F}_{q}^{d}\to\mathbb{R}$. The discrete Laplacian is then
$\Delta g(n)=\sum_{j=1}^{d}\Delta_{j}g(n).$ (4.18)
With these definitions, we will use the following lemma for summation by parts whose proof is elementary and we therefore omit.
Lemma 4.2.
  1. (i)
    For any $g,h:\mathbb{F}_{q}^{d}\to\mathbb{R}$,
    $\sum_{r\in\mathbb{F}_{q}^{d}}\Delta g(r)\,h(r)=\sum_{r\in\mathbb{F}_{q}^{d}}g(r)\,\Delta h(r)$ (4.19)
  2. (ii)
    For every fixed $\ell\in\mathbb{F}_{q}^{d}$,
    $\Delta_{j}\Big{(}e^{-2\pi i\frac{r\cdot\ell}{q}}\Big{)}=-4\sin^{2}\Big{(}\pi\frac{\ell_{j}}{q}\Big{)}\,e^{-2\pi i\frac{r\cdot\ell}{q}},$ (4.20)
    where the derivative is taken in the variable $r\in\mathbb{F}_{q}^{d}$. Therefore,
    $\Delta\Big{(}e^{-2\pi i\frac{r\cdot\ell}{q}}\Big{)}=A(\ell)\,e^{-2\pi i\frac{r\cdot\ell}{q}},\qquad\text{ with }\qquad A(\ell)=-4\sum_{j=1}^{d}\sin^{2}\big{(}\pi\frac{\ell_{j}}{q}\big{)}.$ (4.21)
For $\ell\in\mathbb{F}_{q}^{d}\setminus\{0\}$, from the definition of $\widehat{Z}$ in (4.6) and Lemma 4.2 (ii) we get
$\begin{split}\widehat{Z}(\ell)=\frac{1}{q^{d}}\,\sum_{r\in\mathbb{F}_{q}^{d}}Z(r)\,e^{-2\pi i\frac{r\cdot\ell}{q}}&=\frac{1}{q^{d}\,A(\ell)}\,\sum_{r\in\mathbb{F}_{q}^{d}}Z(r)\,\Delta\Big{(}e^{-2\pi i\frac{r\cdot\ell}{q}}\Big{)}\\ &=\frac{1}{q^{d}\,A(\ell)}\,\sum_{r\in\mathbb{F}_{q}^{d}}\Delta Z(r)\,e^{-2\pi i\frac{r\cdot\ell}{q}},\end{split}$ (4.22)
and iterating,
$\widehat{Z}(\ell)=\frac{1}{q^{d}\,A(\ell)^{\nu}}\,\sum_{r\in\mathbb{F}_{q}^{d}}\Delta^{\nu}Z(r)\,e^{-2\pi i\frac{r\cdot\ell}{q}},\qquad\forall\nu\in\mathbb{N}.$ (4.23)
Hence,
$|\widehat{Z}(\ell)|\leq\frac{1}{q^{d}\,|A(\ell)|^{\nu}}\,\sum_{r\in\mathbb{F}_{q}^{d}}|\Delta^{\nu}Z(r)|,\qquad\forall\nu\in\mathbb{N}.$ (4.24)
We first estimate $\Delta^{\nu}Z(r)$ uniformly in $r$. From the definition of $\zeta$ in (4.3), we get
$|\zeta(n+e_{j})-\zeta(n)|\leq\sup_{x}|\partial_{j}\zeta(x)|\lesssim 1/N+\delta_{j}\lesssim 1/N.$ (4.25)
In the same way,
$|\Delta_{j}\zeta(n)|=\sup_{x}|\partial_{j}(\zeta(x+1)-\zeta(x))|\leq\sup_{x}|\partial_{j}^{2}\zeta(x)|\lesssim\frac{1}{N^{2}},$ (4.26)
and more generally,
$|\Delta_{j_{1}}\ldots\Delta_{j_{\nu}}\zeta(n)|\lesssim\frac{2^{2\nu}}{N^{2\nu}},\qquad\forall\nu\in\mathbb{N},\qquad\forall n\in\mathbb{Z}^{d}.$ (4.27)
Hence, from the definition of $Z$ in (4.5),
$\begin{split}\Delta^{\nu}Z(r)&=\sum_{j_{1}=1}^{d}\cdots\sum_{j_{\nu}=1}^{d}\Delta_{j_{1}}\cdots\Delta_{j_{\nu}}Z(r)\\ &=\sum_{j_{1}=1}^{d}\cdots\sum_{j_{\nu}=1}^{d}\Big{(}\sum_{m\in\mathbb{Z}^{d}}\Delta_{j_{1}}\cdots\Delta_{j_{\nu}}\zeta(mq+r)\Big{)}.\end{split}$ (4.28)
Since $q\ll N$, from the support condition of $\zeta$ we get
$|\Delta^{\nu}Z(r)|\lesssim d^{\nu}\,\Big{(}\frac{N}{q}\Big{)}^{d}\,\Big{(}\frac{2}{N}\Big{)}^{2\nu},\qquad\forall r\in\mathbb{F}_{q}^{d},$ (4.29)
so from (4.24) we get
$|\widehat{Z}(\ell)|\lesssim\Big{(}\frac{N}{q}\Big{)}^{d}\,\Big{(}\frac{4d}{N^{2}|A(\ell)|}\Big{)}^{\nu},\qquad\forall\nu\in\mathbb{N}.$ (4.30)
Joining this with (4.16), we get
$|E|\lesssim(4d)^{\nu}\,q^{d/2}\,\Big{(}\frac{N}{q}\Big{)}^{d}\,\frac{1}{N^{2\nu}}\sum_{\ell\in\mathbb{F}_{q}^{d}\setminus\{0\}}\frac{1}{|A(\ell)|^{\nu}}.$ (4.31)
Calling $\lVert x\rVert=\operatorname{dist}(x,\mathbb{Z})$, we have $|\sin(\pi x)|\simeq\lVert x\rVert$, so
$|A(\ell)|\simeq\sum_{j=1}^{d}\lVert\ell_{j}/q\rVert^{2}.$ (4.32)
This function is symmetric in each component with respect to $[q/2]$, so
$\sum_{\ell\neq 0}\frac{1}{|A(\ell)|^{\nu}}=\sum_{\ell\in\{1,\ldots,[q/2]\}^{d}}\frac{2^{d}}{|A(\ell)|^{\nu}}\simeq 2^{d}q^{2\nu}\sum_{\ell\in\{1,\ldots,[q/2]\}^{d}}\frac{1}{|\ell|^{2\nu}}\simeq C_{d}\,q^{2\nu},$ (4.33)
by choosing any $2\nu>d$, with $C_{d}$ a constant that depends in the dimension $d$ only. Plugging this in (4.31), we get
$|E|\lesssim_{d}q^{d/2}\,\Big{(}\frac{N}{q}\Big{)}^{d}\,\Big{(}\frac{q}{N}\Big{)}^{2\nu}.$ (4.34)
Estimate for the solution
In view of the choice of $x$ and $t$ in (4.1) and of the estimate (4.15) holding for $b\in G(q)$, we define
$X_{N}=\bigcup_{q\in\mathcal{P}_{Q}}\,\bigcup_{b\in G(q)}B\Big{(}\frac{b}{q},\frac{1}{N}\Big{)},\qquad\text{ where }\qquad\mathcal{P}_{Q}=\{\,q\simeq Q\,:\,q\text{ prime }\,\}.$ (4.35)
Thus, if $x\in X_{N}$, we have $x\in B(b/q,1/N)$ for some $q\in\mathcal{P}_{Q}$ and $b\in G(q)$, so choosing $t=1/q$ both estimates (4.15) and (4.34) hold. Since $Q\ll N$, we conclude that for large enough $N\gg 1$,
$\forall x\in X_{N},\quad\exists t\simeq\frac{1}{Q}\quad\text{ such that }\quad\Big{|}e^{itP(D)}f_{N}(x)\Big{|}\gtrsim\Big{(}\frac{N}{\sqrt{Q}}\Big{)}^{d}.$ (4.36)
4.2. The set of divergence
To compute the measure of the set $X_{N}$ defined in (4.35), the basic idea boils down to controling the overlap among the balls in the union. For a collection of sets $\{E_{j}\}_{j\in J}$, define $f=\sum_{j}\chi_{E_{j}}$. Then,
$\Big{(}\sum_{j}|E_{j}|\Big{)}^{2}=\lVert f\rVert_{L^{1}}^{2}\leq\operatorname{supp}(f)\,\lVert f\rVert_{L^{2}}^{2}\leq\Big{|}\bigcup_{j}E_{j}\Big{|}\,\sum_{j,k}|E_{j}\cap E_{k}|$ (4.37)
The number of pairs of sets $(E_{j},E_{k})$ that overlap is at least $|J|$, because each set overlaps itself. Suppose that there is not much more overlapping, that is, that the number of pairs $(E_{j},E_{k})$ that overlap is bounded by $C|J|$ for some $C>0$. Then, if the sets have all comparable measures, say $|E_{j}|\simeq M$, from (4.37) we get
$M|J|\sum_{j}|E_{j}|\lesssim\Big{(}\sum_{j}|E_{j}|\Big{)}^{2}\leq\Big{|}\bigcup_{j}E_{j}\Big{|}\,\sum_{j,k}|E_{j}\cap E_{k}|\lesssim\Big{|}\bigcup_{j}E_{j}\Big{|}\,C\,|J|\,M.$ (4.38)
This implies
$\sum_{j}|E_{j}|\lesssim\Big{|}\bigcup_{j}E_{j}\Big{|},$ (4.39)
with means that, in terms of measure, the sets $\{E_{j}\}$ are basically disjoint. In the case of $X_{N}$ in (4.35), the total number of balls in the union is
$|J|=\sum_{q\in\mathcal{P}_{Q}}\sum_{b\in G(q)}1\simeq\frac{Q^{d+1}}{\log Q}.$ (4.40)
On the other hand, the condition for two balls to overlap is
$\Big{|}\frac{b_{i}}{q}-\frac{b^{\prime}_{i}}{q^{\prime}}\Big{|}\leq\frac{2}{N},\quad\forall i=1,\ldots,d,\qquad\Longleftrightarrow\qquad|b_{i}q^{\prime}-b_{i}^{\prime}q|\leq\frac{2\,q\,q^{\prime}}{N},\quad\forall i=1,\ldots,d.$ (4.41)
To count the number of such pairs, we recall the following lemma from elementary number theory. We include its proof in Appendix A for completeness.
Lemma 4.3.
Let $q,q^{\prime}\in\mathbb{N}$. The number of integer pairs $(b,b^{\prime})$ that satisfy $1\leq b\leq q$ and $1\leq b^{\prime}\leq q^{\prime}$ and that
$0<|bq^{\prime}-b^{\prime}q|\leq A$ (4.42)
is bounded by $2A$.
With this lemma in hand, fix $q,q^{\prime}\in\mathcal{P}_{Q}$. If $q=q^{\prime}$, given that $Q\ll N$, the condition (4.41) holds only when $b_{i}=b_{i}^{\prime}$ for all $i$, so there are #$\#G(q)\simeq q^{d}$ such pairs $(b,b^{\prime})$. If $q\neq q^{\prime}$, by Lemma 4.3 we have at most $(4qq^{\prime}/N)^{d}$ such pairs $(b,b^{\prime})$. In all, thus, the number of intersecting pairs is at most
$\sum_{q\in\mathcal{P}_{Q}}q^{d}+\sum_{q\neq q^{\prime}\in\mathcal{P}_{Q}}\Big{(}\frac{4qq^{\prime}}{N}\Big{)}^{d}\simeq Q^{d}\frac{Q}{\log Q}+\Big{(}\frac{Q^{2}}{N}\Big{)}^{d}\Big{(}\frac{Q}{\log Q}\Big{)}^{2}\simeq\frac{Q^{d+1}}{\log Q}\,\Big{(}1+\frac{Q^{d+1}}{N^{d}\log Q}\Big{)}.$ (4.43)
Since $Q^{d+1}=N^{d}$, this is bounded by $Q^{d+1}/\log Q\simeq|J|$. Hence, the discussion leading to (4.38) implies that
$|X_{N}|\gtrsim\sum_{q\in\mathcal{P}_{Q}}\sum_{b\in G(q)}\Big{|}B\Big{(}\frac{b}{q},\frac{1}{N}\Big{)}\Big{|}\simeq\frac{Q^{d+1}}{\log Q}\,\frac{1}{N^{d}}\simeq\frac{1}{\log Q}.$ (4.44)
4.3. Disproving the maximal estimate
Gathering it all, from (4.36) we have the following for every $N\gg 1$:
$\forall x\in X_{N},\quad\exists t\simeq N^{-d/(d+1)}\quad\text{ such that }\quad\Big{|}e^{itP(D)}f_{N}(x)\Big{|}\gtrsim\Big{(}\frac{N}{\sqrt{Q}}\Big{)}^{d}.$ (4.45)
Joining this with the computation of $\lVert f_{N}\rVert_{L^{2}}$ in (2.3) and of $|X_{N}|$ in (4.44), we get
$\frac{\big{\lVert}\,\sup_{t}\big{|}e^{itP(D)}f_{N}\big{|}\,\big{\rVert}_{L^{2}}}{\lVert f_{N}\rVert_{L^{2}}}\gtrsim|X_{N}|^{1/2}\,\Big{(}\frac{N}{\sqrt{Q}}\Big{)}^{d}\,N^{-d/2}\simeq\frac{1}{(\log N)^{1/2}}\,N^{\frac{d}{2(d+1)}}.$ (4.46)
Hence, if $s<\frac{d}{2(d+1)}$, we get
$\frac{\big{\lVert}\,\sup_{t}\big{|}e^{itP(D)}f_{N}\big{|}\,\big{\rVert}_{L^{2}}}{\lVert f_{N}\rVert_{H^{s}}}\gtrsim\frac{1}{(\log N)P^{1/2}}\,N^{\frac{d}{2(d+1)}-s}\to\infty,\qquad\text{ as }N\to\infty,$ (4.47)
and the maximal estimate is disproved.
Appendix A Proof of Lemma 4.3
For the sake of completeness, we provide a geometrically motivated proof of Lemma 4.3. Having $A>0$ and $q,q^{\prime}\in\mathbb{N}$ fixed, and rewriting inequality (4.42) as
$0<\frac{|bq^{\prime}-b^{\prime}q|}{\sqrt{q^{2}+(q^{\prime})^{2}}}\leq\frac{A}{\sqrt{q^{2}+(q^{\prime})^{2}}},$ (A.1)
we can interpret a pair $(b,b^{\prime})$ satisfying (A.1) as a lattice point whose distance to the line $L$ given by the equation $q^{\prime}x-qy=0$ is less than $A/\sqrt{q^{2}+(q^{\prime})^{2}}$ (see Figure 1). In this setting, it suffices to bound the number of lattice points in the diagonal rectangle shown in Figure 1 by $2A$.
Since $b,b^{\prime},q,q^{\prime}\in\mathbb{N}$, these lattice points satisfy $bq^{\prime}-b^{\prime}q=i$ for some $i\in\mathbb{Z}$ with $0<|i|\leq A$. Moreover, denoting $d=\operatorname{gcd}(q,q^{\prime})$, we directly get $d\mid i$. Hence,
$bq^{\prime}-b^{\prime}q=k\,d,\qquad\text{ with }\quad k\in\mathbb{Z},\quad 0<|k|\leq\lfloor A/d\rfloor,$ (A.2)
which represents a line parallel to $L$ that we denote by $L_{k}$. There are at most $2A/d$ such lines. Hence, it suffices to bound the number of lattice points in each $L_{k}$ by $d$. To do that, suppose that $(b,b^{\prime})\in L_{k}$ is the first lattice point with $b>0$. If we write $q=d\,m$ and $q^{\prime}=d\,m^{\prime}$ such that $\operatorname{gcd}(m,m^{\prime})=1$, then
$b^{\prime}=\frac{m^{\prime}}{m}\,b-\frac{k}{m}\in\mathbb{Z}.$ (A.3)
Then, for other lattice points $(b+\beta,b^{\prime}+\beta^{\prime})\in\mathbb{Z}^{2}$ with $\beta\geq 0$ we have
$(b+\beta,b^{\prime}+\beta^{\prime})\in\mathbb{Z}^{2}\cap L_{k}\quad\Longrightarrow\quad\frac{m^{\prime}}{m}\,(b+\beta)-\frac{k}{m}\in\mathbb{Z}\quad\Longrightarrow\quad\frac{m^{\prime}}{m}\,\beta\in\mathbb{Z},$ (A.4)
and therefore $m\mid\beta$. Hence, the abcissa component of the lattice points in $L_{k}$ have $1\leq b+\kappa\,m\leq q$ with $0\leq\kappa<q/m=d$. Consequently, each $L_{k}$ has at most $d$ lattice points, and since there are at most $2A/d$ lines $L_{k}$, we conclude that the maximum number of lattice points in the region is $d\times 2A/d=2A$. ∎
Figure 1. Lemma 4.3 counts the number of lattice points in the intersection of the diagonal rectangle with $[1,q]\times[1,q^{\prime}]$ that do not lie on the diagonal line $L$.
Acknowledgements
We would like to thank Chu-Hee Cho and Lillian Pierce for clarifying discussions about the topic.
This project has received funding from the European Union’s Horizon Europe research and innovation programme under Marie Sklodowska Curie Actions with grant agreement 101104250 - TIDE. D.E. acknowledges the support of the Simons Foundation Collaboration Grant on Wave Turbulence (Nahmod’s award ID 651469) and the American Mathematical Society and the Simons Foundation under an AMS-Simons Travel Grant for the period 2022-2024. X.Y. is partially supported by NSF DMS-2306429.